solve each system of equatioins by using substitution
m+n=20
m-n=-4
work: m+n=20
-n -n
----------
20-n
m-n=-4
20-n-n=-4
(20-n)=-4
-4n-80
~ thats all i could come up with....

Answer by checkley77(12844)

m+n=20
m=20-n
Now replace m in the following equation with 20-n
m-n=-4
(20-n)-n=-4
-2n=-4-20
-2n=-24
n=-24/-2
n=12 ans.
m+12=20
m=20-12
m=8 ans.